OpenStudy (anonymous):

Find the missing term. 7^-147 × 7^98 = 7^18 × 7^-38 × A)7^(-20) B)7^22 C)7^(-29) D)7^58

OpenStudy (tkhunny):

Have you considered adding exponents and seeing what's missing?

OpenStudy (anonymous):

yes @tkhunny

OpenStudy (tkhunny):

Okay, what were the results of that effort?

OpenStudy (anonymous):

I did not find the answer :( I do not understand @tkhunny

OpenStudy (tkhunny):

-147 + 98 = ??

OpenStudy (anonymous):

-49? @tkhunny

OpenStudy (anonymous):

When multiplying powers with the same base, the result is simplified by adding the exponents. For example\[3^{-25}\times 3^{14} = 3^{-25+14} = 3^{-11}\]So, on order for this equation to be true, the sum of the exponents on both sides of the equation must be equal. Sum the exponents on the left hand side and make the sum of the exponents on the right hand the same.

OpenStudy (tkhunny):

18 - 38 + What = -49